02:04 LBCB Trackers failure

02:26 Boot successful

02:00 Around this time, DX was not raising - finally did at about 02:30

02:49 Test flats (took data but telescope not ready yet -just to set up filter/rotator)

02:54 Test flats - this time ready, but OB was in 'play' mode yet showing no status updates. DX had panicked...that was why.

03:00 Closing because RH was rising - SMT is now 91.9, LBT says 94.3 at 20:01

08:55 Calling it a night. The RH has been high and for the period when it did drop below 90%, there were thick clouds & no stars visible. Later, ~08:00, the WinCC? (?) computer which controls temperature & pressures for the HBS was not working. It continuously rebooted itself. Calls were made but this was not resolved. Urgent issuetrak filed because we will need this for Sunday night.

-- MarcoPedani? - 24 May 2009
I Attachment Action Size Date Who Comment
daq.20090524.gifgif daq.20090524.gif manage 49 K 17 Jul 2009 - 05:03 JohnHill Plot of telescope and mirror temperatures
Topic revision: r2 - 17 Jul 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback